One kilogram of water at 5˚C is heated to 100˚C. Compute its change in entropy. Assume that the specific heat of water is constant at 4190 J/kg·K over this temperature range

1
Expert's answer
2021-11-25T10:36:54-0500

The change in entropy is


"ΔS=2.303nC_P\log\frac{T_2}{T_1},\\\space\\
ΔS=2.303\frac m{\Mu}C_P\log\frac{T_2}{T_1},\\\space\\=68.4\text{ kJ}."